# Enhancing Ethereum Fraud Detection via Generative and Contrastive Self-supervision

### v1

@@ V-D Evaluation on Fraud Detection
To evaluate the effectiveness and superiority of our Meta-IFD, we compare it against 10 baselines on two fraud datasets. The detection results are reported in Table II, from which we can draw the following conclusions:
[…]
@@ V-E1 Impact of View Setting
As shown in Table III, when our Meta-IFD is adapted to a single-view setting, i.e., Meta-IFD (single view), its performance in detecting Ponzi schemes and phishing scams decreases by 1.71% and 0.91% (Macro-F1 metric), respectively, indicating that the utilization of multi-view settings can enhance the characterization of account features to some extent. Furthermore, under the multi-view setting, we vary the number of views within {2, 3, 4, 5, 10, 20}, as depicted in Fig. 7, from which we can observe that: 1) As the number of views varies, the Macro-F1 metric fluctuates more significantly compared to Micro-F1, suggesting that the former is more effective in evaluating framework performance under imbalanced data scenarios; 2) The framework’s performance (measured by Macro-F1 metric) shows a general trend of initially increasing and then decreasing with the number of views. This suggests that increasing the number of views appropriately is beneficial for the detection of minority class fraud samples, but too many views may lead to feature redundancy, which in turn interferes with analysis of complex Ethereum interaction behaviors.
TABLE III: Performance comparison between Meta-IFD and its ablation models.
@@
Meta-IFD (single view) 91.97$\pm$0.69 80.56$\pm$2.19 97.36$\pm$0.71 95.29$\pm$2.25
Meta-IFD (weight sharing) 91.08$\pm$0.67 79.05$\pm$1.54 96.46$\pm$0.18 93.06$\pm$0.28
@@ V-E2 Impact of Weight Separation Setting
To validate the effectiveness of the account type-specific transformation in our Meta-IFD, we compare the performance of the framework under a weight-sharing setting, i.e., Meta-IFD (weight sharing), as shown in Table III. The results show that weight separation learning for different types of account features yields positive gains, suggesting an enhanced distinguishability among the features of distinct accounts.
@@ V-E3 Impact of Self-attention Setting
To validate the superiority of the self-attention in inter-view feature aggregation, we compare three aggregation operations: summation, concatenation, and naive attention. As shown in Table III, the self-attention aggregation operation facilitates our Meta-IFD in achieving improvement ranging from 2.02% to 3.89% and 0.87% to 2.78% on the two datasets (Macro-F1 metric), respectively. This result demonstrates that the self-attention mechanism can more effectively extract and fuse key features from different views, thereby enhancing fraud detection.
Figure 8: Illustration of the meta-interaction feature distance margin distribution before and after introducing contrastive self-supervision.
@@
V-F Analysis of Coarse-grained Feature Contract Module
@@ V-F1 Effect of Contrastive Self-supervision
As shown in Table III, the removal of the contrast module results in a performance degradation of 11.65% and 1.81% on the two datasets, respectively. This observation highlights the effective improvement of our framework’s performance in fraud detection through feature contrast. Additionally, we compare the distance margin distribution between different types of meta-interaction features before and after introducing contrastive self-supervision, as shown in Fig. 8. It can be observed that after introducing contrastive self-supervision, the peak of the distance margin distribution shifts significantly to the right, indicating that the distance between anchors $r^{2}_{\ast}$ and positive samples $r^{3}_{\ast}$ is reduced, while the distance between anchors and negative samples $r^{1}_{\ast}$ is increased. This phenomenon further illustrates that introducing contrastive self-supervision facilitates Meta-IFD in learning the similarities among interactions of the same type, as well as distinctions between interactions of different types, improving its ability to distinguish different interaction behaviors and ultimately enhancing fraud detection.
@@ V-F2 Impact of Trade-off Coefficient
Furthermore, we investigate how the framework’s performance is affected by the intensity of the auxiliary module, which is controlled by the trade-off coefficient $\lambda$. We vary $\lambda$ within {0.001, 0.01, 0.1, 1, 10, 100} and observe the fluctuations in the framework’s performance, as shown in Fig. 9. Firstly, we can see that in detecting Ponzi schemes, Meta-IFD achieves better detection results when $\lambda$ is larger. Combined with the data statistics from Table I and the distance margin distribution in Fig. 8, a reasonable explanation is that the Ponzi scheme dataset is more severely affected by data distribution imbalance and label scarcity issues, and higher intensity of contrast self-supervised constraint are needed to help Meta-IFD better distinguish different interaction behavior patterns. Therefore, appropriately increasing the $\lambda$ setting helps improve the framework’s detection performance. For the phishing scam dataset, its data distribution is relatively balanced, and the label size is relatively abundant compared to the Ponzi scheme dataset, which facilitates effective learning of account behavior representations within Meta-IFD. Consequently, only moderate contrastive self-supervision constraint is required to enhance the discriminative capability of our framework.

### v2

@@ V-D Evaluation on Fraud Detection
To evaluate the effectiveness and superiority of our Meta-IFD, we compare it against 12 baselines on two fraud datasets. The detection results are reported in Table IV, from which we can draw the following conclusions.
Overall, compared with the optimal machine learning, homogeneous graph analysis, heterogeneous graph analysis methods, and sequence analysis-based methods, our Meta-IFD achieves 14.77%, 4.05%, 6.76%, and 0.77% improvement in detecting Ponzi schemes, and 2.02%, 0.83%, 0.95%, and 1.15% in detecting phishing scams (Macro-F1 metric), which demonstrates the effectiveness and superiority of our Meta-IFD in Ethereum fraud detection.
To evaluate the performance in detecting fraudulent accounts, Table IV presents the precision, recall, and F1-score under binary evaluation metrics. As observed, our Meta-IFD achieves the best performance in terms of F1-score and ranks second in both precision and recall. Taking Ponzi detection as an example, Trans2Vec performs well in precision but poorly in recall, indicating that this method is overly conservative in detecting fraud, leading to a high rate of missed detections. Conversely, BERT4ETH performs well in recall but poorly in precision, suggesting that this approach tends to misclassify normal accounts as fraudulent, resulting in a high false positive rate. The F1-score balances these two aspects, and the best F1 performance demonstrates that our Meta-IFD provides a more stable performance in fraud detection.
The detection performance of different types of baselines varies significantly across various fraud categories. We select representative baselines for analysis. LightGBM achieves relatively powerful performance in phishing detection but performs poorly in Ponzi scheme detection, indicating that the general manual features defined by machine learning methods cannot adequately describe diverse fraudulent behaviors. To achieve better detection performance, machine learning methods necessitate meticulous design of manual features tailored to different fraud categories, which lacks good generalization and heavily depends on expert knowledge. The two-stage homogeneous graph analysis methods represented by Trans2Vec suffer from similar limitations, in addition to the fact that it focuses more on the structural information of the transaction graph. Other major research in this field face similar challenges. For example, BERT4ETH, a more recent advancement that incorporates Transformer models for fraud detection, demonstrates high performance on Ponzi detection but exhibits poor generalizability when detecting diverse fraudulent behaviors. The ZipZap method, an improvement based on BERT4ETH, achieves some enhancement in generalizability but at the cost of reduced performance in certain aspects. This indicates that achieving both high generalizability and high effectiveness in Ethereum fraud detection remains a significant technical challenge. Other end-to-end heterogeneous graph analysis methods show moderate detection performance for different fraudulent behaviors, somewhat surpassing homogeneous graph analysis methods. A reasonable explanation is that both are affected by the interaction graph construction and data sampling methods. Finally, our Meta-IFD achieves the best detection performance for different frauds, indicating its generalizability and effectiveness in fraud detection.
@@ V-E1 Impact of View Setting
As shown in Table V-E1, when our Meta-IFD is adapted to a single-view setting, i.e., “Module Removal ($w/o\ M$)”, its performance in detecting Ponzi schemes and phishing scams decreases by 4.89% and 0.34% (Macro-F1 metric), respectively, indicating that the utilization of multi-view settings can enhance the characterization of account features to some extent. Furthermore, under the multi-view setting, we vary the number of views within {1, 2, 3, 4, 5, 10, 20} , as depicted in Fig. 7, from which we can observe that: The framework’s performance (measured by Macro-F1 metric) shows a general trend of initially increasing and then decreasing with the number of views. This suggests that increasing the number of views appropriately is beneficial for the detection of minority class fraud samples, but too many views may lead to feature redundancy, which fails to improve the detection performance and even in turn interferes with analysis of complex Ethereum interaction behaviors.
TABLE V: Performance comparison between Meta-IFD and its ablation models.
@@
Module Removal $w/o\ I$ 91.67$\pm$0.87 80.13$\pm$1.82 97.61$\pm$0.45 97.58$\pm$0.45
$w/o\ M$ 90.48$\pm$1.07 78.69$\pm$1.29 97.61$\pm$0.39 97.57$\pm$0.39
$w/o\ \mathcal{L}_{\text{con}}$ 91.23$\pm$0.73 80.42$\pm$1.33 97.76$\pm$0.56 97.72$\pm$0.56
@@ V-E2 Impact of Weight Separation Setting
To validate the effectiveness of the account type-specific transformation in our Meta-IFD, we compare the performance of the framework under a weight-sharing setting, i.e., “Others (Sharing)”, as shown in Table V-E1. The results show that weight separation learning for different types of account features yields positive gains, suggesting an enhanced distinguishability among the features of distinct accounts.
@@ V-E3 Impact of Self-attention Setting
To validate the superiority of the self-attention in inter-view feature aggregation, we compare three aggregation operations: summation, concatenation, and naive attention. As shown in Table V-E1, the self-attention aggregation operation facilitates our Meta-IFD in achieving improvement ranging from 1.59% to 3.71% and 0.22% to 0.33% on the two datasets (Macro-F1 metric), respectively. This result demonstrates that the self-attention mechanism can more effectively extract and fuse key features from different views, thereby enhancing fraud detection.
Figure 8: Illustration of the meta-interaction feature distance margin distribution before and after introducing contrastive self-supervision. The bars denote the distribution of the number of accounts, with their width signifying the size of the statistical window. Curves represent trends in the distribution.
@@
V-F Analysis of Coarse-grained Feature Contrast Module
@@ V-F1 Effect of Contrastive Self-supervision
As shown in Table V-E1, the removal of the contrast module results in a performance degradation of 2.64% and 0.18% on the two datasets, respectively. This observation highlights the effective improvement of our framework’s performance in fraud detection through feature contrast. Additionally, we compare the distance margin distribution between different types of meta-interaction features before and after introducing contrastive self-supervision, as shown in Fig. 8. We utilize the Euclidean distance to quantify the similarity between samples, assessing the discrepancy in similarity scores between positive and negative samples. Subsequently, we record both the count of accounts falling within the specified similarity window and their corresponding trend. It can be observed that after introducing contrastive self-supervision, the peak of the distance margin distribution shifts significantly to the right, indicating that the distance between anchors $r^{2}_{\ast}$ and positive samples $r^{3}_{\ast}$ is reduced, while the distance between anchors and negative samples $r^{1}_{\ast}$ is increased. This phenomenon further illustrates that introducing contrastive self-supervision facilitates Meta-IFD in learning the similarities among interactions of the same type, as well as distinctions between interactions of different types, improving its ability to distinguish different interaction behaviors and ultimately enhancing fraud detection.
@@ V-F2 Impact of Trade-off Coefficient
Furthermore, we investigate how the framework’s performance is affected by the intensity of the auxiliary module, which is controlled by the trade-off coefficient $\lambda$. We vary $\lambda$ within {0.001, 0.01, 0.1, 1, 10, 100} and observe the fluctuations in the framework’s performance, as shown in Fig. 9. Firstly, we can see that in detecting Ponzi schemes, Meta-IFD achieves better detection results when $\lambda$ is larger. Combined with the data statistics from Table III and the distance margin distribution in Fig. 8, a reasonable explanation is that the Ponzi scheme dataset is more severely affected by data distribution imbalance and label scarcity issues, and higher intensity of contrast self-supervised constraint are needed to help Meta-IFD better distinguish different interaction behavior patterns. Therefore, appropriately increasing the $\lambda$ setting helps improve the framework’s detection performance. For the phishing scam dataset, its data distribution is relatively balanced, and the label size is relatively abundant compared to the Ponzi scheme dataset, which facilitates effective learning of account behavior representations within Meta-IFD. Consequently, only moderate contrastive self-supervision constraint is required to enhance the discriminative capability of our framework.
V-G Analysis of Feature Generation
V-G1 Impact of Input Features
To validate the effectiveness and superiority of the handcrafted features proposed in this work, we separately extract features based on call and trans interactions, i.e., we retain only the 7-dimensional features derived from a single interaction type for subsequent fraud detection. For fairness, we extend these features to the same 14 dimensions. The experimental results are shown in Table V-E1 under “Feature (call)” and “Feature (trans)”. It can be observed that the Macro-F1 score drops by 6.74% and 1.40%, respectively. Notably, detection using “Feature (trans)” outperforms that using “Feature (call)”, particularly in phishing detection. This is because phishing fraud typically occurs between EOAs, involving fewer call behaviors. Furthermore, trans, as a more common behavior, carries richer semantic information. In summary, considering different interaction behaviors during feature construction enables the acquisition of a more comprehensive feature set, thereby improving the effectiveness of subsequent fraud detection.
V-G2 Impact of Using Interaction Types as Conditions in Feature Generation
To validate the effectiveness of incorporating interaction types as conditions during the feature generation process, we remove the interaction-aware component, degrading ICVAE into a standard CVAE model for interaction feature generation. The fraud detection performance of this ablation model is presented in Table. V-E1 under “Module Removal $(w/o\ I)$”. The results show that the macro-F1 scores drop by 3.01% and 0.33%, respectively. Notably, detecting Ponzi schemes is more dependent on interaction information, as Ponzi schemes often exploit more interaction behaviors to achieve their fraudulent objectives. Therefore, incorporating interaction information during the pretraining phase effectively captures fine-grained behavioral patterns of accounts. In conclusion, integrating interaction types as conditions during the feature generation stage allows for the creation of richer features for low-degree nodes, thereby improving subsequent detection performance.
Figure 10: Performance and computational efficiency of different methods. Colors indicate different methods and sizes indicate Macro-F1 values.
V-H Analysis of Computational Efficiency
We further investigate the computational efficiency of different methods. Considering that the fine-grained interaction feature generation module can be pre-trained, we focus on evaluating the computational efficiency of the end-to-end multi-view feature learning stage. We compare only end-to-end detection methods under similar settings. Fig. 10 reports the total runtime and the number of epochs required for convergence under early stopping for each method. For fairness, we set the hidden layer size to 64, repeat experiments five times, and report the average results. We can observe that while Meta-IFD incurs slightly higher total runtime compared to some methods, it converges in fewer epochs and achieves superior performance. This indicates that our Meta-IFD is capable of learning data features more efficiently. Additionally, when we remove the contrastive learning module—a time-intensive component—we find that the runtime decreased significantly. However, the number of epochs required for convergence increases substantially (particularly on the phishing dataset), suggesting that the self-supervised module improves the efficiency of feature learning. The results also highlights that heterogeneous methods generally require higher time costs than homogeneous methods. This is because heterogeneous methods must differentiate between node and edge types, leading to increased computational complexity.